# supplement

### v1

@@ 5 Iterative HEAT Estimation
First-Round Communication:
  1. 1.
    Sample Splitting:
    For $m\in[M]$, at the $m$-th site, do:
    1. (a)
      Randomly split the samples into two disjoint sets $\mathcal{I}^{(m_{1})}$ and $\mathcal{I}^{(m_{2})}$ such that $[n_{m}]=\mathcal{I}^{(m_{1})}\bigcup\mathcal{I}^{(m_{2})}$ and $|\mathcal{I}^{(m_{2})}|=\kappa_{m}n_{m}$ for some constant $\kappa_{m}\in(0,1)$.
    2. (b)
      Compute $\widehat{\Omega}^{(m_{1})}$ in (21) based on $\{X_{i}^{(m)},i\in\mathcal{I}^{(m_{1})}\}$, and construct $\widehat{\Sigma}^{(m_{2})}$ based on $\{X_{i}^{(m)},i\in\mathcal{I}^{(m_{2})}\}$.
    3. (c)
      Transmit the proportion $\kappa_{m}$ to the central site.
  2. 2.
    Initialization:
    Implement Algorithm 3, and obtain the first-round HEAT estimator $\widetilde{\Omega}^{(m)}$ for $m\in[M]$. Define $\widetilde{\Omega}^{(m,1)}:=\widetilde{\Omega}^{(m)}$, $m\in[M]$.
Multi-Round Communications:For the iterations $t=2,\ldots,T$, do:
@@
      Symmetrized debiased estimation:
      $\widebar{\Omega}^{(m,t)}=\frac{1}{2}\left(\widebar{\Omega}^{(m,t-1/2)}+(\widebar{\Omega}^{(m,t-1/2)})^{\scriptscriptstyle\sf T}\right),$
      where $\widebar{\Omega}^{(m,t-1/2)}=(\widehat{\Omega}^{(m_{1})})^{\scriptscriptstyle\sf T}+\widetilde{\Omega}^{(m,t-1)}-(\widehat{\Omega}^{(m_{1})})^{\scriptscriptstyle\sf T}\widehat{\Sigma}^{(m_{2})}\widetilde{\Omega}^{(m,t-1)}$.
@@
      $\widetilde{\Omega}^{(m,t)}=\widehat{\Gamma}^{(t)}+\widehat{\Lambda}^{(m,t)}=(\widehat{\Gamma}_{j,k}^{(t)})_{j,k\in[p]}+(\widehat{\Lambda}_{j,k}^{(m,t)})_{j,k\in[p]},$
@@
      $\widehat{\Gamma}_{j,k}^{(t)}=T_{1,\lambda_{1,j,k}^{(t)}}(\widebar{\Omega}_{j,k}^{(t)}),\widehat{\Lambda}_{j,k}^{(\cdot,t)}=T_{2,\lambda_{2,j,k}^{(t)}}(\widebar{\Omega}_{j,k}^{(\cdot,t)}-\widebar{\Omega}_{j,k}^{(t)}1_{M}),$
@@ 5.1 IteHEAT Algorithm
IteHEAT begins with a preliminary data splitting step and initialization using Algorithm 3, and then refines the estimates through iterative local debiasing and central integration. This approach offers insights into the statistical refinement of higher-order errors and the accuracy enhancement in integrative analysis.
First, at each site, the dataset is randomly split into two disjoint subsets, and then these two sets are separately used to construct the regularized estimation for $\Omega^{(m)}$ and the covariance estimation for $\Sigma^{(m)}$. Specifically, for each $m\in[M]$, the $m$-th site randomly splits the samples into two disjoint sets, $\mathcal{I}_{1}^{(m_{1})}$ and $\mathcal{I}^{(m_{2})}$, such that $[n_{m}]=\mathcal{I}^{(m_{1})}\bigcup\mathcal{I}^{(m_{2})}$ and $|\mathcal{I}^{(m_{2})}|=\kappa_{m}n_{m}$, where the splitting proportion $\kappa_{m}\in(0,1)$ is a preassigned universal constant. The superscripts $m_{1}$ and $m_{2}$ indicate the first and second subsets at the $m$-th site, respectively. For $a\in\{1,2\}$, we define the subsets $X^{(m_{a})}:=X_{\mathcal{I}^{(m_{a})},[p]}^{(m)}\in\mathbb{R}^{|\mathcal{I}^{(m_{a})}|\times p}$ and the centralized subsets $\widecheck{X}^{(m_{a})}=(\widecheck{X}^{(m_{a})}_{i,j})\in\mathbb{R}^{|\mathcal{I}^{(m_{a})}|\times p}$, where $\widecheck{X}^{(m_{a})}_{i,j}=X^{(m_{a})}_{i,j}-\widebar{X}^{(m_{a})}_{j}$ and $\widebar{X}^{(m_{a})}=\frac{1}{|\mathcal{I}^{(m_{a})}|}\sum_{i\in\mathcal{I}^{(m_{a})}}X_{i}^{(m)}$. Similar to the local estimator $\widehat{\Omega}^{(m)}$ in (4)-(5), the regularized estimation $\widehat{\Omega}^{(m_{1})}$ is constructed based on the first subset $X^{(m_{1})}$:
$\widehat{\Omega}_{j,j}^{(m_{1})}=\frac{(1-\kappa_{m})n_{m}}{(\widecheck{X}_{\cdot,j}^{(m_{1})})^{\scriptscriptstyle\sf T}(\widecheck{X}_{\cdot,j}^{(m_{1})}-\widecheck{X}_{\cdot,-j}^{(m_{1})}\widehat{\gamma}_{j}^{(m_{1})})},\text{ and }\widehat{\Omega}_{-j,j}^{(m_{1})}=-\widehat{\gamma}_{j}^{(m_{1})}\widehat{\Omega}_{j,j}^{(m_{1})},$ (21)
@@
$\widehat{\gamma}_{j}^{(m_{1})}=\arg\min_{\gamma\in\mathbb{R}^{p-1}}\left\{\frac{1}{2(1-\kappa_{m})n_{m}}\|\widecheck{X}_{\cdot,j}^{(m_{1})}-\widecheck{X}_{\cdot,-j}^{(m_{1})}\gamma\|_{2}^{2}+\frac{\lambda_{j}^{(m)}}{\sqrt{1-\kappa_{m}}}\left\|\gamma\right\|_{1}\right\}.$ (22)
Besides, the empirical covariance $\widehat{\Sigma}^{(m_{2})}:=\frac{1}{|\mathcal{I}^{(m_{2})}|}(\widecheck{X}^{(m_{2})})^{\scriptscriptstyle\sf T}\widecheck{X}^{(m_{2})}$ is constructed based on the second subsets. For $m\in[M]$, the proportion $\kappa_{m}$ is sent to the central site.
Next, Algorithm 3 is implemented to obtain the initial first-round HEAT estimator $\widetilde{\Omega}^{(\cdot,1)}:=\widetilde{\Omega}^{(\cdot)}$, followed by multi-round communications to achieve the IteHEAT estimation. Specifically, for $t\geq 2$, during the $t$-th round of communication, the $m$-th site computes the symmetrized debiased estimator by
@@
$\widebar{\Omega}^{(m,t-1/2)}=(\widehat{\Omega}^{(m_{1})})^{\scriptscriptstyle\sf T}+\widetilde{\Omega}^{(m,t-1)}-(\widehat{\Omega}^{(m_{1})})^{\scriptscriptstyle\sf T}\widehat{\Sigma}^{(m_{2})}\widetilde{\Omega}^{(m,t-1)}.$ (24)
@@
$\widetilde{\Omega}^{(m,t)}=\widehat{\Gamma}^{(t)}+\widehat{\Lambda}^{(m,t)}=(\widehat{\Gamma}_{j,k}^{(t)})_{j,k\in[p]}+(\widehat{\Lambda}_{j,k}^{(m,t)})_{j,k\in[p]},$ (25)
@@
$\widehat{\Gamma}_{j,k}^{(t)}=T_{1,\lambda_{1,j,k}^{(t)}}(\widebar{\Omega}_{j,k}^{(t)}),\text{ and }\widehat{\Lambda}_{j,k}^{(\cdot,t)}=T_{2,\lambda_{2,j,k}^{(t)}}(\widebar{\Omega}_{j,k}^{(\cdot,t)}-\widebar{\Omega}_{j,k}^{(t)}1_{M}),$
and $\widebar{\Omega}^{(t)}=\frac{1}{N}\sum_{m=1}^{M}n_{m}\widebar{\Omega}^{(m,t)}$. Note that, the entry-adaptive shrinkage levels $\lambda_{1,j,k}^{(t)},\lambda_{2,j,k}^{(t)}\geq 0$ at the $t$-th iteration depend on the splitting proportions and the variance estimation (9), and will be specified in (26)-(27) of the subsequent section.
@@
The data splitting strategy is employed in the debiasing steps (24) of IteHEAT. Despite the efficiency loss induced by splitting, the independence structure between $\widehat{\Omega}^{(m_{1})}$ and $\widehat{\Sigma}^{(m_{2})}$ facilitates the technical analysis, and similar approaches are utilized in Cai et al., (2023); Guo, (2023) for statistical inference. Numerically, the simulation studies in Section 6 and Section LABEL:sec:splitting_effect of the supplement demonstrate the superior performance of IteHEAT without splitting. This finding is consistent with Cai et al., (2023); Guo, (2023) and suggests that non-split IteHEAT is preferable for practical implementation.
@@ 5.2 Theoretical Properties
Next, we establish the error analysis for the IteHEAT estimators $\widetilde{\Omega}^{(\cdot,t)},t\geq 2$ under loss matrices (17)-(18). The shrinkage levels $\lambda_{1,j,k}^{(t)},\lambda_{2,j,k}^{(t)},t\geq 2$ in (25) are specified as
$\lambda_{1,j,k}^{(t)}=(2+\delta)\sqrt{A_{j,k}\frac{\log(p)}{N}}+\widetilde{C}_{1}\left(\frac{M\log(p)}{N}+s_{0}\sqrt{\frac{\log(p)}{n}}(\max_{j,k\in[p]}\lambda_{1,j,k}^{(t-1)}+\max_{j,k\in[p]}\lambda_{2,j,k}^{(t-1)})\right),$ (26)
$\lambda_{2,j,k}^{(t)}=$ $(1+\delta)\sqrt{\frac{B_{1,j,k}+2\sqrt{2}B_{2,j,k}\sqrt{\log(p)}+4B_{\infty,j,k}\log(p)}{N}}$ (27)
$+\widetilde{C}_{2}\left(\frac{\sqrt{M}\log(p)}{\sqrt{nN}}+s_{0}\sqrt{\frac{\log(p)}{n}}(\max_{j,k\in[p]}\lambda_{1,j,k}^{(t-1)}+\max_{j,k\in[p]}\lambda_{2,j,k}^{(t-1)})\right),$
with $\lambda_{1,j,k}^{(1)}:=\lambda_{1,j,k}$ and $\lambda_{2,j,k}^{(1)}:=\lambda_{2,j,k}$, where $A_{j,k}=\sum_{m\in[M]}\frac{n_{m}}{\kappa_{m}N}\widehat{v}_{j,k}^{(m)}$, $B_{1,j,k}=\sum_{m=1}^{M}\widehat{v}_{j,k}^{(m)}/\kappa_{m}$, $B_{2,j,k}=\sqrt{\sum_{m=1}^{M}(\widehat{v}_{j,k}^{(m)}/\kappa_{m})^{2}}$, $B_{\infty,j,k}=\max_{m\in[M]}\widehat{v}_{j,k}^{(m)}/\kappa_{m}$, $\delta>0$ is given in Equations (19)-(20), and $\widetilde{C}_{1},\widetilde{C}_{2}>0$ are some universal constants independent of $(n,p,t)$. The following theorems present iterative contraction rates of convergence for IteHEAT.
@@
Under Assumptions 1-5, if $n\geq Cs_{0}^{2}\log(p)$ for some constant $C>0$, $s_{0}\log(p)\log^{2}(p\vee n_{\max})=o(n)$, $M^{4}\log^{4}(p\vee n_{\max})=o(n)$ and $\lambda_{j}^{(m)}\asymp\sqrt{\log(p)/n_{m}}$ uniformly for $j\in[p]$ and $m\in[M]$, then for $t\geq 2$, $r\geq 1$ and $l\in[1,\infty]$, we have
@@
and $\varphi_{1}^{\star}=\sqrt{\frac{\log(p)}{N}}+\frac{M\log(p)}{N}$, $\varphi_{2}^{\star}=\sqrt{\frac{M+\log(p)}{N}}+\frac{\sqrt{M}\log(p)}{\sqrt{nN}}$.
@@
The results of IteHEAT are analogous to those of HEAT in Theorems 1-2, with $\varphi_{1}$ and $\varphi_{2}$ replaced by $\varphi_{1}^{(t)}$ and $\varphi_{2}^{(t)}$, respectively. Both $\varphi_{1}^{(t)}$ and $\varphi_{2}^{(t)}$ consist of iteration-independent and iteration-dependent terms, which respectively represent the leading errors and the iteratively refined higher-order errors. Notably, the additional terms $\frac{M\log(p)}{N}$ and $\frac{\sqrt{M}\log(p)}{\sqrt{nN}}$ in $\varphi_{1}^{\star}$ and $\varphi_{2}^{\star}$ are induced by the mean estimation for $\mu^{(m)},m\in[M]$. They are included here for generality and will vanish if the means are known (e.g., $\mu^{(m)}=0$) or if the means are assumed to be homogeneous across distributed sites and are estimated using the full datasets.
The factor $s_{0}\sqrt{\log(p)/n}$ in the iteration-dependent terms serves as the contraction rate for higher-order error reduction and leads to the geometric convergence. As a result, the number of iterations $T$ with the logarithm complexity
$T\gtrsim 1+\log\left(\frac{\sqrt{M+\log(p)}+s_{0}\sqrt{M/n}\log(p)}{\sqrt{\log(p)}+M/\sqrt{N}\log(p)}\right)/\log\left(\frac{1}{s_{0}}\sqrt{\frac{n}{\log(p)}}\right)$
@@
$\varphi_{1}^{(T)}\asymp\varphi_{1}^{\star}+s_{0}\sqrt{\frac{\log(p)}{n}}\varphi_{2}^{\star},\text{ and }\varphi_{2}^{(T)}\asymp\varphi_{2}^{\star}.$
That is, the higher-order errors induced by aggregating local estimators are dominated by the leading errors after logarithm rounds of communication between the central and local sites. Note that, the contraction results indicate that the local sample sizes cannot be excessively small, a requirement also noted in iterative distributed algorithms under homogeneity (e.g., Wang et al.,, 2017; Jordan et al.,, 2019; Chen et al.,, 2019). Consequently, with a few rounds of iterations, the results in Theorems 5-6 match the minimax lower bounds in Theorems 3-4 if $\frac{M^{2}}{N}\lesssim(\log(p))^{-1}$, $\log^{2}(p)\lesssim(1\vee\frac{\log(p)}{M})n$ and $M\vee\log(p)\lesssim ns_{0}^{-2}$. Therefore, IteHEAT enhances accuracy and optimality compared to HEAT by leveraging iterative refinements.

### v2

@@ 5 Iterative HEAT Estimation
2 Iterative Distributed Estimation for Heterogeneous Precision Matrices
First-Round Communication: 1. Sample Splitting: For $m\in[M]$, at the $m$-th site, do: (a) Randomly split the samples into $H$ disjoint sets $\mathcal{I}_{1}^{(m)},\cdots,\mathcal{I}_{H}^{(m)}$ such that $n_{m}]=\bigcup_{h\in[H]}\mathcal{I}_{h}^{(m)}$ and $|\mathcal{I}_{h}^{(m)}|\asymp n_{m}/H$. (b) For $h\in[H]$, compute $\widehat{\Omega}_{-h}^{(m)}$ in (22) based on $\{X_{i}^{(m)},i\in\mathcal{I}_{-h}^{(m)}:=\bigcup_{h^{\prime}\neq h}\mathcal{I}_{h^{\prime}}^{(m)}\}$. 2. Initialization: Implement Algorithm 3, and obtain the first-round HEAT estimator $\widetilde{\Omega}^{(m)}$ for $m\in[M]$. Define $\widetilde{\Omega}^{(m,1)}:=\widetilde{\Omega}^{(m)}$, $m\in[M]$. Multi-Round Communications:For the iterations $t=2,\ldots,T$, do:
@@ 2 Iterative Distributed Estimation for Heterogeneous Precision Matrices
      Cross-fitted and aggregated debiasing: $\widebar{\Omega}^{(m,t-1/2)}=n_{m}^{-1}\sum_{h\in[H]}|\mathcal{I}_{h}^{(m)}|\widebar{\Omega}_{h}^{(m,t-1/2)}$, where
      $\widebar{\Omega}_{h}^{(m,t-1/2)}=(\widehat{\Omega}_{-h}^{(m)})^{\scriptscriptstyle\sf T}+\widetilde{\Omega}^{(m,t-1)}-(\widehat{\Omega}_{-h}^{(m)})^{\scriptscriptstyle\sf T}\widehat{\Sigma}_{h}^{(m)}\widetilde{\Omega}^{(m,t-1)},$
      and $\widehat{\Sigma}_{h}^{(m)}$ is the empirical covariance matrix based on $\{X_{i}^{(m)},i\in\mathcal{I}_{h}^{(m)}\}$.
@@
      Symmetrization: $\widebar{\Omega}^{(m,t)}=\frac{1}{2}\left(\widebar{\Omega}^{(m,t-1/2)}+(\widebar{\Omega}^{(m,t-1/2)})^{\scriptscriptstyle\sf T}\right).$
    3. (c)
@@
      $\widetilde{\Omega}^{(m,t)}=\widetilde{\Gamma}^{(t)}+\widetilde{\Lambda}^{(m,t)}=(\widetilde{\Gamma}_{j,k}^{(t)})_{j,k\in[p]}+(\widetilde{\Lambda}_{j,k}^{(m,t)})_{j,k\in[p]},$
@@
      $\widetilde{\Gamma}_{j,k}^{(t)}=T_{1,\lambda_{1,j,k}^{(t)}}(\widebar{\Omega}_{j,k}^{(t)}),~\widetilde{\Lambda}_{j,k}^{(\cdot,t)}=T_{2,\lambda_{2,j,k}^{(t)}}(\widebar{\Omega}_{j,k}^{(\cdot,t)}-\widebar{\Omega}_{j,k}^{(t)}1_{M}),$
@@ 5.1 IteHEAT Algorithm
IteHEAT begins with a preliminary data splitting and initialization using Algorithm 3, and then refines the estimates through iterative local debiasing and central integration. This approach offers insights into the statistical refinement of higher-order errors and the accuracy enhancement in integrative analysis.
First, for a fixed integer $H\geq 2$, each site splits the data into $H$ disjoint subsets, and these subsets are used to construct $H$ regularized estimators for $\Omega^{(m)}$. Specifically, for each $m\in[M]$, the $m$-th site randomly splits the samples into $H$ disjoint folds $\mathcal{I}_{1}^{(m)},\cdots,\mathcal{I}_{H}^{(m)}$ such that $n_{m}]=\bigcup_{h\in[H]}\mathcal{I}_{h}^{(m)}$ and $|\mathcal{I}_{h}^{(m)}|\asymp n_{m}/H$, and we denote $\mathcal{I}_{-h}^{(m)}:=\bigcup_{h^{\prime}\neq h}\mathcal{I}_{h^{\prime}}^{(m)}$. Based on this partition, we define the data $X_{h}^{(m)}:=X_{\mathcal{I}_{h}^{(m)},\cdot}^{(m)}\in\mathbb{R}^{|\mathcal{I}_{h}^{(m)}|\times p}$ and $X_{-h}^{(m)}:=X_{\mathcal{I}_{-h}^{(m)},\cdot}^{(m)}\in\mathbb{R}^{(n_{m}-|\mathcal{I}_{h}^{(m)}|)\times p}$ as well as the corresponding centralized data $\widecheck{X}_{h}^{(m)}\in\mathbb{R}^{|\mathcal{I}_{h}^{(m)}|\times p}$ and $\widecheck{X}_{-h}^{(m)}\in\mathbb{R}^{(n_{m}-|\mathcal{I}_{h}^{(m)}|)\times p}$, where $(\widecheck{X}_{h}^{(m)})_{i,j}=(X_{h}^{(m)})_{i,j}-(\widebar{X}_{h}^{(m)})_{j}$ and $(\widecheck{X}_{-h}^{(m)})_{i,j}=(X_{-h}^{(m)})_{i,j}-(\widebar{X}_{-h}^{(m)})_{j}$ and empirical means $\widebar{X}_{h}^{(m)}=\frac{1}{|\mathcal{I}_{h}^{(m)}|}\sum_{i\in\mathcal{I}_{h}^{(m)}}X_{i}^{(m)}$ and $\widebar{X}_{-h}^{(m)}=\frac{1}{|\mathcal{I}_{-h}^{(m)}|}\sum_{i\in\mathcal{I}_{-h}^{(m)}}X_{i}^{(m)}$. Then, similar to the local estimator $\widehat{\Omega}^{(m)}$ in (4)-(5), for each fold $h\in[H]$, the regularized estimation $\widehat{\Omega}_{-h}^{(m)}\in\mathbb{R}^{p\times p}$ is constructed based on the data $X_{-h}^{(m)}$:
[…]
@@
$\widehat{\gamma}_{-h,j}^{(m)}=\arg\min_{\gamma\in\mathbb{R}^{p-1}}\left\{\frac{1}{2|\mathcal{I}_{-h}^{(m)}|}\|(\widecheck{X}_{-h}^{(m)})_{\cdot,j}-(\widecheck{X}_{-h}^{(m)})_{\cdot,-j}\gamma\|_{2}^{2}+\sqrt{\frac{n_{m}}{|\mathcal{I}_{-h}^{(m)}|}}\lambda_{j}^{(m)}\left\|\gamma\right\|_{1}\right\}.$ (23)
Thus, in contrast to the estimator $\widehat{\Omega}^{(m)}$ in Algorithm 3, the $\widehat{\Omega}_{-h}^{(m)}$ is obtained by excluding the samples in subset $\mathcal{I}_{h}^{(m)}$.
Next, Algorithm 3 is implemented to obtain the initial first-round HEAT estimator $\widetilde{\Omega}^{(\cdot,1)}:=\widetilde{\Omega}^{(\cdot)}$, followed by multi-round communications to achieve the IteHEAT estimation. For $t\geq 2$, during the $t$-th round of communication, the $m$-th site computes the symmetrized debiased estimator by
@@
$\widebar{\Omega}^{(m,t-1/2)}=\sum_{h\in[H]}\frac{|\mathcal{I}_{h}^{(m)}|}{n_{m}}\widebar{\Omega}_{h}^{(m,t-1/2)}$ (25)
is the aggregated estimator across the multiple split folds and
$\widebar{\Omega}_{h}^{(m,t-1/2)}=(\widehat{\Omega}_{-h}^{(m)})^{\scriptscriptstyle\sf T}+\widetilde{\Omega}^{(m,t-1)}-(\widehat{\Omega}_{-h}^{(m)})^{\scriptscriptstyle\sf T}\widehat{\Sigma}_{h}^{(m)}\widetilde{\Omega}^{(m,t-1)},~h\in[H].$ (26)
Here, $\widehat{\Sigma}_{h}^{(m)}$ denotes the empirical covariance matrix based on the subset $\mathcal{I}_{h}^{(m)}$, i.e., $\widehat{\Sigma}_{h}^{(m)}=\frac{1}{|\mathcal{I}_{h}^{(m)}|}(\widecheck{X}_{h}^{(m)})^{\scriptscriptstyle\sf T}\widecheck{X}_{h}^{(m)}$. The debiasing steps in (25) and (26) mainly incorporate the idea of cross-fitting (Chernozhukov et al.,, 2018, 2022) into the construction in (6), and the step in (24) ensures the matrix symmetrization.
@@
$\widetilde{\Omega}^{(m,t)}=\widetilde{\Gamma}^{(t)}+\widetilde{\Lambda}^{(m,t)}=(\widetilde{\Gamma}_{j,k}^{(t)})_{j,k\in[p]}+(\widetilde{\Lambda}_{j,k}^{(m,t)})_{j,k\in[p]},$ (27)
@@
$\widetilde{\Gamma}_{j,k}^{(t)}=T_{1,\lambda_{1,j,k}^{(t)}}(\widebar{\Omega}_{j,k}^{(t)}),\text{ and }\widetilde{\Lambda}_{j,k}^{(\cdot,t)}=T_{2,\lambda_{2,j,k}^{(t)}}(\widebar{\Omega}_{j,k}^{(\cdot,t)}-\widebar{\Omega}_{j,k}^{(t)}1_{M}),$
and $\widebar{\Omega}^{(t)}=\frac{1}{N}\sum_{m=1}^{M}n_{m}\widebar{\Omega}^{(m,t)}$. Similar to the determination of $\lambda_{1,j,k}$ and $\lambda_{2,j,k}$ as detailed in Section 3.2 and Section 4.1, either universal shrinkage or entry-dependent adaptive shrinkage levels can be applied for $\lambda_{1,j,k}^{(t)}$ and $\lambda_{2,j,k}^{(t)}$ at the $t$-th iteration.
@@
In spite of the independence between $\widehat{\Sigma}_{h}^{(m)}$ and $\widehat{\Omega}_{-h}^{(m)}$, the debiasing step in (26) is not a standard cross-fitting procedure as in Chernozhukov et al., (2018, 2022), and the distinction arises because the IteHEAT estimator $\widetilde{\Omega}^{(m,t-1)}$ from the previous (t1)(t-1)-th round is dependent with the full local data $X^{(m)}$. Nevertheless, this procedure yields valid debiased estimator in (24)-(25), and the inclusion of $\widetilde{\Omega}^{(m,t-1)}$ serves to iteratively refine the higher-order errors of the local estimators.
@@ 5.2 Theoretical Properties
Next, we establish the error analysis for the IteHEAT estimators $\widetilde{\Omega}^{(\cdot,t)},t\geq 2$ in (27) under loss matrices (17)-(18). For the shrinkage levels $\lambda_{1,j,k}^{(t)}$ and $\lambda_{2,j,k}^{(t)}$ specified in Equations (LABEL:eq:lambda_1jk_2jk_univ_ite)-(LABEL:eq:lambda_1jk_2jk_adap_ite) in Section LABEL:sec:theorem_56 of the supplement, the following theorems present iterative contraction rates of convergence for IteHEAT.
@@
Under Assumptions 1-5, if $s_{0}\log(p)=o(n)$, $\log^{3}(p)=o(N)$, $n\geq Cs_{0}^{2}\log(p)$ for some constant $C>0$, $\tau^{2}=o(M\vee\log(p))$ and $\lambda_{j}^{(m)}\asymp\sqrt{\log(p)/n_{m}}$ uniformly for $j\in[p]$ and $m\in[M]$, and additionally assume $\log^{2}(p\vee n_{\max})(s_{0}\vee\log(p))=o(n)$ for the shrinkage levels in Equation (LABEL:eq:lambda_1jk_2jk_adap_ite), then for $t\geq 2$, $r\geq 1$ and $l\in[1,\infty]$, we have
@@
and $\varphi_{1}^{\star}=\sqrt{\frac{\log(p)}{N}}+\frac{M}{N}$, $\varphi_{2}^{\star}=\sqrt{\frac{M+\log(p)}{N}}+\sqrt{\frac{\log(p)}{N}\frac{s_{0}\log^{2}(p)}{n}}$.
@@
The results for IteHEAT are analogous to those for HEAT in Theorems 1-2, with $\varphi_{1}$ and $\varphi_{2}$ replaced by $\varphi_{1}^{(t)}$ and $\varphi_{2}^{(t)}$, respectively. For the iteration-dependent errors in $\varphi_{1}^{(t)}$ and $\varphi_{2}^{(t)}$, the factor $s_{0}\sqrt{\log(p)/n}$ serves as the contraction rate for iterative higher-order error reduction and $\left(s_{0}\sqrt{\log(p)/n}\right)^{t-1}$ leads to the geometric convergence. Such contraction implies that condition $n\gtrsim s_{0}^{2}\log(p)$ is required for convergence, i.e., the local sample sizes cannot be excessively small. Similar requirements are also present in other iterative distributed algorithms designed under homogeneity (Wang et al.,, 2017; Chen et al.,, 2019; Jordan et al.,, 2019; Tan et al.,, 2022). As a result, the number of iterations $T$ with the logarithm complexity
$T\gtrsim 1+\log\left(\frac{\sqrt{M+\log(p)}+s_{0}\sqrt{M/n}\log(p)}{\sqrt{\log(p)}+M/\sqrt{N}}\right)/\log\left(\frac{1}{s_{0}}\sqrt{\frac{n}{\log(p)}}\right)$
@@
$\varphi_{1}^{(T)}\asymp\varphi_{1}^{\star}+s_{0}\sqrt{\frac{\log(p)}{n}}\varphi_{2}^{\star},\text{ and }\varphi_{2}^{(T)}\asymp\varphi_{2}^{\star}.$ (28)
In other words, the iterative refined higher-order errors induced by aggregating local estimators are dominated by the fixed statistical errors after logarithm rounds of communication between the central and local sites.
Notably, the convergence rates in (28) match the optimal rates established in Theorems 3-4, up to additional terms $\frac{M}{N}$, $\sqrt{\frac{\log(p)}{N}\frac{s_{0}\log^{2}(p)}{n}}$ and $s_{0}\sqrt{\frac{\log(p)}{n}}\varphi_{2}^{\star}$. The two terms $\frac{M}{N}$ and $\sqrt{\frac{\log(p)}{N}\frac{s_{0}\log^{2}(p)}{n}}$ arise from estimating the individual means $\widebar{X}^{(m)},m\in[M]$ for data centralization, which relies only on the local samples. They get dominated by $\psi_{1}$ and $\psi_{2}$ under the mild conditions $M\lesssim\sqrt{N\log(p)}$ (or $M\lesssim n\log(p)$) and $s_{0}\log^{2}(p)\lesssim n$. Moreover, these data centralization errors would vanish if the means are known (e.g., $\mu^{(m)}=0$ for $m\in[M]$), or become negligible if the means are assumed to be homogeneous across distributed sites and are estimated using the pooled empirical means.
The remaining error $s_{0}\sqrt{\frac{\log(p)}{n}}\varphi_{2}^{\star}$ represents the cost for estimating homogeneous components via iterative aggregation in the presence of heterogeneity. It is dominated by the leading error $\psi_{1}$ if $\frac{\sqrt{M}}{n}\lesssim(s_{0}^{2}\log(p))^{-1}$ and $M\lesssim\frac{n}{s_{0}^{2}}\vee\log(p)$, which are weaker than those required for HEAT as shown in Section 4.1. It is also worth noting that, this error term would vanish in homogeneous setting where $s_{2}=0$, which highlights the distinction between distributed estimation under homogeneity versus heterogeneity.
Consequently, after a few rounds of communication, the results in Theorems 5-6 show that IteHEAT enhances the estimation accuracy and optimality compared to HEAT by leveraging iterative refinements.